The power of a lens is expressed in diopters (D).
Answer #1


combined power P = P1 + P2 + P3 = 1D + 2D + 3D = 6D <----answer

P1 = 1/f1


f1 = 1/P1 = 1/1 = 1m = 100cm <---answer

for second lens

P2 = 1/f2


f2 = 1/p2 = 1/2 = 0.5m = 50 cm <--answer


for third lens


P3 = 1/f3

f3 = 1/p3 = 1/3 = 0.333 m = 33.3 cm <-answer

combined focal length 1/F = 1/f1 + 1/f2 + 1/f3

1/F = P1 + P2 + P3 = P


1/F = P


F = 1/P = 1/6 m


focal length of the stack = F = 100 /6 cm = 16.7 cm <------answer